Show that the function omega = sin z satisfies the Cauchy-Riemann and Laplace
equations
1
Expert's answer
2020-03-05T11:09:45-0500

The Cauchy-Riemann equation:


Let "f(x,y)=u(x,y)+iv(x,y)" , where "u" and "v" are real.

Then u=sinxcoshy and v=cosxsinhy st.


"u_x=\cos x\cosh y = v_y,~~~~~~~v_x=-\sin x\sinh y = - u_y,"


i.e., CR conditions hold.


"f(z)=\frac{(e^{ix}-e^{-ix})(e^y+e^{-y})}{4i}+i\frac{(e^{ix}+e^{-ix})(e^y-e^{-y})}{4}=\frac{e^{ix}e^{-y}-e^{-ix}e^{y}}{2i}= \frac{e^{i(x+iy)}-e^{-i(x+iy)}}{2i}=\sin z;"


The Laplace equation:


"\Omega_{x}=\cos(x+iy),\Omega_{xx}=-\sin(x+iy),"


"\Omega_{y}=\cos (x+iy)*i, \Omega_{yy}=-\sin(x+iy)*i^2=\sin(x+iy)."


Thus, "\Omega_{xx}+\Omega_{yy}=0"